OpenStudy (anonymous):

State the phase shift of the function. y=10sin(theta over 4 -4pi) -5

OpenStudy (anonymous):

4pi right

OpenStudy (anonymous):

thank you, but wouldn't you have to divide the 4pi by 1/4? @ECE

OpenStudy (anonymous):

Probably, depends what kind of eqn you're using

OpenStudy (anonymous):

does it really say \[\sin(\frac{\theta}{4}-4\pi)\]?

OpenStudy (anonymous):

if it does, then there is no phase shift, because sine is periodic with period \(2\pi\) and so \[\sin(\frac{\theta}{4}-4\pi)=\sin(\frac{\theta}{4})\]

OpenStudy (anonymous):

If it's in the form \[Asin(\omega*t + \theta) \] then yea 4pi right

OpenStudy (anonymous):

i guess what i am trying to say is that there is no such thing as a phase shift of \(4\pi\) or any integer multiple of \(2\pi\)

OpenStudy (anonymous):

yes @satellite73 it does say that

OpenStudy (anonymous):

ok then there is no phase shift

OpenStudy (anonymous):

i mean it says y=10sin(theta over 4-4pi)-5 @satellite73